OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation. x + 6y + z = 2 x – 3y + 2z = –1 x + 5y + 3z = –4

OpenStudy (anonymous):

Solve [1] for x: x =-6y-z+2 Now sub it into [2] in place of x: (-6y-z+2)-3y+2z=-1 -9y+z=-3 [2]' Now sub it into [3] in place of x: (-6y-z+2)+5y+3z=-4 -y+2z=-6 [3]' The new 2x2 system would then be solved: -9y+z=-3 [2]' -y+2z=-6 [3]' Your problem implies that this is the answer, not the actual solution to the 3x3 system.

OpenStudy (anonymous):

thank you for your help..but are you sure??

OpenStudy (anonymous):

well i have made arithmetic errors before; let me check it with wolfram

OpenStudy (anonymous):

im gonna go with your answer...thnkx

OpenStudy (anonymous):

the answer to the 3x3 is x=5, y=0, z=-3; the answer to the 2x2 is y=0 z=-3 yes I am sure